OpenStudy (anonymous):

What is the sum of the infinite geometric series? mc011-1.jpg mc011-2.jpg mc011-3.jpg –4 –6

OpenStudy (anonymous):

We can't see what the series are. However, for a general geometric series, you have \[\sum_{n=0}^\infty ar^n=\begin{cases}\dfrac{a}{1-r}&\text{for }|r|<1\\\\ \pm\infty&\text{otherwise}\end{cases}\]